How many grams of oxygen are needed to combust 20.0 grams of propane?
1
Expert's answer
2020-04-03T11:52:35-0400

propane = C3H8.

M(C3H8) = 3*Ar(C) + 8*Ar(H) = 3*12 + 8*1 = 44 (g/mol).


Solution:

The equation of a chemical reaction:

C3H8 + 5O2 = 3CO2 + 4H2O.

According to the chemical equation: n(C3H8) = n(O2)/5.

n(O2) = 5 * n(C3H8).


n(C3H8) = m(C3H8) / M(C3H8) = (20.0 g) / (44 g/mol) = 0.4545 moles.

Then,

n(O2) = 5 * n(C3H8) = 5 * 0.4545 = 2.2725 moles.

m(O2) = n(O2) * M(O2) = (2.2725 moles) * (32 g/mol) = 72.72 g.

m(O2) = 72.72 g.


Answer: 72.72 grams of oxygen (O2) are needed to combust 20.0 grams of propane.